OpenStudy (anonymous):

How many solutions will this system of equations have? y = 3.5x - 3.5 y = -3.5x + 3.5 No solution Infinite solutions One solution Two solutions Medal and Fan

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

I got it its b

OpenStudy (anonymous):

can you answer the question iclosed idk it

OpenStudy (anonymous):

Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5

OpenStudy (texaschic101):

same slope, different y intercepts -- no solutions because they are parallel. same slope, same y intercept -- infinite solutions because they are the same line different slopes, different y intercepts -- 1 solution Do you know how to find the slopes and y intercepts of these equations ?

OpenStudy (anonymous):

its the 3.5x

OpenStudy (anonymous):

Idk how to do anything on the second question

OpenStudy (texaschic101):

y = mx + b form...the slope is in the m position and the y intercept is in the b position. y = 3.5x - 3.5 --- slope is 3.5 and y int is -3.5 y = -3.5x + 3.5 -- slope is -3.5 and y intercept is 3.5 the slopes are different and the y intercepts are different (because of the signs). Therefore, there is only one solution

OpenStudy (anonymous):

so its one solution? and whats the second question this one Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5

OpenStudy (texaschic101):

I did this one earlier...but it confuses me because I do not see how a set of equations can be inconsistent and independent. I do know that 2x + y = -6 would cause an inconsistent equation because the slopes of the lines are the same and the y intercepts are different...meaning parallel lines and inconsistent

OpenStudy (anonymous):

so its 2x+y=-6?

OpenStudy (texaschic101):

That would cause an inconsistent set of equations...so I think so